OpenStudy (anonymous):

limx->0 sin(abs(x))/x If I am correct, than it is undefined? I mean... It isn't differentiable at all values because abs(x) isn't differentiable? So I can't do L'Hospitals?? @Mathematics

OpenStudy (anonymous):

\[\lim_{x\rightarrow 0}\frac{\sin(|x|)}{x}\]?

OpenStudy (anonymous):

think of it like this: Case 1: x > 0 (so x approaches 0 from above): = limx->0 sin(x)/x = 1 Case 2: x < 0 (so x approaches 0 from below): = limx->0 (sin(-x))/x = limx->0 -sin(x)/x = -1 therefore the limit does not exist

OpenStudy (anonymous):

yeah satellite

OpenStudy (anonymous):

are you sure that limx->0+ of sin(x)/x =1??

OpenStudy (anonymous):

jamesm nailed it. on one direction you get 1, in the other -1, no limit

OpenStudy (anonymous):

it is always the case that \[\lim_{x\rightarrow 0}\frac{\sin(ax)}{bx}=\frac{a}{b}\]

OpenStudy (anonymous):

Does that work with all trig functions?

OpenStudy (anonymous):

so \[\lim_{x\rightarrow 0}\frac{\sin(x)}{x}=1\] and \[\lim_{x\rightarrow 0}\frac{\sin(-x)}{x}=-1\]

OpenStudy (anonymous):

And I can't apply L'hospital's to that sucker either right..?

OpenStudy (anonymous):

yes, it's true that limx->0+ of sin(x)/x = 1 more generally sin(x)/x goes to 1 as x approaches 0 from both sides

OpenStudy (anonymous):

no because it is not differentiable at 0

OpenStudy (anonymous):

Okay, well what about this? \[\lim_{x \rightarrow 0} (x-2)^{2}\sin(\pi/x-5) \]

OpenStudy (anonymous):

\[l \lim_{x \rightarrow 5} (x-2)^{2} =9 \] and \[\lim_{x \rightarrow 5} \sin(\pi/x-5) = \sin(0) = 1\] Which makes 1/0... So what do I do now?

OpenStudy (anonymous):

limit as 0 or at 5?

OpenStudy (anonymous):

this line \[\lim_{x \rightarrow 5} \sin(\pi/x-5) = \sin(0) = 1\] is not correct.

myininaya (myininaya):

pi/(x-5) or pi/x-5?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

Oh crap. Okay so sin(pi/0) =??? No idea what to do!

OpenStudy (anonymous):

division by zero is undefined

OpenStudy (anonymous):

\[\lim_{x\rightarrow 5}\sin(\frac{\pi}{x-5})\] does not exist

myininaya (myininaya):

\[\lim_{x \rightarrow 0} (x-2)^{2}\sin(\frac{\pi}{x-5}) \text{ is this the problem } \]

OpenStudy (anonymous):

So the answer to this is undefined? WolframAlpha is saying "Undefined on the interval (-9,9)". Is this what is expected of me on a quiz? where did they get the -9,9.

OpenStudy (anonymous):

No myininaya, the limit of x is approaching 5

OpenStudy (anonymous):

as x goes to 5, x - 5 goes to zero and \[\frac{\pi}{x-5}\] goes to infinity. the sine therefore takes on all values between -1 and 1 infinitely often. there is no limit here

OpenStudy (anonymous):

But wait! we have to multiply everything by 9. AHA, i get it!

OpenStudy (anonymous):

Sweet, okay that makes sense. That's how WolframA got the interval (-9,9). Thanks so much! that was a tricky one for me..

OpenStudy (anonymous):

to be more precise \[\frac{\pi}{x-5}\] goes to infinity if x > 5 and minus infinity if x < 5 in any case there is no limit. the 9 is a red herring and has nothing to do with the problem

OpenStudy (anonymous):

Okay, I'm glad you mentioned that last part. So I would not mention "on the interval (-9,9)" on a quiz. Thanks!

OpenStudy (anonymous):

So I would say on a quiz or something: lim x->5 of sin(pi/x-5) is undefined, therefore \[\lim_{x \rightarrow 5}(5-2)^{2}\sin(\pi/(x-5)\] is undefined as x is approaching 5?